# Equitable Federated Learning with Activation Clustering

### v1

@@ 2 Related Work
Clustering and Fairness in Federated Learning. Preserving privacy by not sharing data or sensitive client information with the server or other clients is a crucial aspect of FL, posing challenges for developing learning algorithms. Some work [15, 16, 17] prioritize client selection to boost performance using local validation losses. We demonstrate client discrepancies using the global model’s performance at the end of each epoch, though this information is not essential for our algorithm’s efficiency. Other approaches [18, 19] group the clients based on their representations using similarity matrices or submodular sets. These methods assume a fixed number of clients per round, deviating from FL standards. Approaches like [20, 21] prioritize highly contributing clients, undermining the consensus-based nature of FL. The method in [22] minimizes the maximum loss across all data samples to avoid bias towards any data distribution, differing from our server-side debiasing approach using activation vector information. Additionally, [23] proposes domain adaptation for client groups with similar characteristics, whereas we focus on disjoint client groups participating in the collaboration. Other works such as [24] aim to mitigate the bias between the data samples with a client’s data to promote group fairness, whereas we focus on fairness among groups of clients where the goal is to create equitable learning scenarios among different groups. A recent study [25] also focuses on ensuring fair performance for both groups of clients and individual clients. However, prior knowledge of the client groups is required, while our approach automatically clusters the clients based on the activation vectors. Another concurrent work [26] orthogonal to ours uses the bias’s gradients in the neural network’s last layer to construct clusters and then perform heterogeneity-aware client sampling. This work poses an immense computational overhead, requiring computing the similarity matrix amongst all clients to sample the participating clients. Other clustering-based works [27, 28] perform clustered federated learning to promote personalization, which is not the objective of our work.
@@ 4.2 Framework: Equitable-FL
We emphasize that while K-Means relies on specifying the number of clusters as a hyperparameter, it is not unique in this requirement. Many other clustering algorithms also depend on a thresholding mechanism to determine cluster formation, which similarly requires fine-tuning for optimal results. Thus, the need for parameter adjustment is a common aspect across clustering techniques.
@@ 4.3 Convergence Analysis
For $\eta_{k}LE\leq\frac{1}{2}$, $\eta_{k}\mu E\leq\frac{1}{2}$, $\mu<1$, and let Assumptions 1, 2, 3, and 4 hold true for Equitable-FL (refer Algorithm 1). In Equitable-FL, set $\eta_{k}=\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)$ for all $k$. Define a distribution $\mathbb{P}$ for $k\in\{0,\ldots,K-1\}$ such that $\mathbb{P}(k)=\frac{(1+\zeta-1)^{(K-1-k)}}{\sum_{k=0}^{K-1}(1+\zeta_{1})^{k}}$ where $\zeta:=\eta_{k}^{2}E^{2}\left(9\eta_{k}L^{2}E+4\eta_{k}\mu E+6L\left(1+\frac{4\eta_{k}^{2}\mu^{2}E^{2}}{18}\right)\right)$. Sample $k^{*}$ from $\mathbb{P}$ uniformly. Then:
$\mathbb{E}[\|\nabla f(\bm{w}_{k^{*}})\|^{2}]=\mathcal{O}\left(\frac{1}{\sqrt{K}}\sigma^{2}\right).$ (8)
@@ 6 Conclusion
Limitation and Future works: As previously mentioned, privacy is a cornerstone of Federated Learning (FL). However, activation vectors can lead to information leakage if the server is compromised. Despite this risk, privacy-preserving mechanisms exist to enable secure learning without compromising privacy [40, 41]. This paper aims to demonstrate the effectiveness of using activation vectors to cluster client groups and develop a fair weighting mechanism for these groups. We also propose that this clustering approach can be applied to personalized federated learning [42, 43]. This method effectively clusters client groups and ensures fair solutions for each group, maintaining fairness for all participants.
@@ Appendix A Theorems and Lemmas
Theorem 2.
(Detailed proof of Theorem 1) For $\eta_{k}LE\leq\frac{1}{2}$, $\eta_{k}\mu E\leq\frac{1}{2}$, $\mu<1$, and let Assumptions 1, 2, 3, and 4 hold true for Equitable-FL (refer Algorithm 1). In Equitable-FL, set $\eta_{k}=\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)$ for all $k$. Define a distribution $\mathbb{P}$ for $k\in\{0,\ldots,K-1\}$ such that $\mathbb{P}(k)=\frac{(1+\zeta-1)^{(K-1-k)}}{\sum_{k=0}^{K-1}(1+\zeta_{1})^{k}}$ where $\zeta:=\eta_{k}^{2}E^{2}\left(9\eta_{k}L^{2}E+4\eta_{k}\mu E+6L\left(1+\frac{4\eta_{k}^{2}\mu^{2}E^{2}}{18}\right)\right)$. Sample $k^{*}$ from $\mathbb{P}$ uniformly. Then:
$\mathbb{E}[\|\nabla f(\bm{w}_{k^{*}})\|^{2}]=\mathcal{O}\left(\frac{1}{\sqrt{K}}\sigma^{2}\right)$
@@
Now, note that the optimal step size will be $\eta=\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)$. Then
$\sum_{k=0}^{K-1}p_{k}\mathbb{E}[\|\nabla f(\bm{w}_{k})\|^{2}]=\mathcal{O}\left(\frac{1}{\sqrt{K}}\sigma^{2}\right)$ (17)

### v2

@@ 2 Related Work
Clustering and Fairness in Federated Learning. Preserving privacy by not sharing data or sensitive client information with the server or other clients is a crucial aspect of FL, posing challenges for developing learning algorithms. Some work [15, 16, 17] prioritize client selection to boost performance using local validation losses. We demonstrate client discrepancies using the global model’s performance at the end of each epoch, though this information is not essential for our algorithm’s efficiency. Other approaches [18, 19, 20] group the clients based on their representations using similarity matrices or submodular sets. These methods assume a fixed number of clients per round, deviating from FL standards. Approaches like [21, 22] prioritize highly contributing clients, undermining the consensus-based nature of FL. The method in [23] minimizes the maximum loss across all data samples to avoid bias towards any data distribution, differing from our server-side debiasing approach using activation vector information. Additionally, [24] proposes domain adaptation for client groups with similar characteristics, whereas we focus on disjoint client groups participating in the collaboration. Other works such as [25] aim to mitigate the bias between the data samples with a client’s data to promote group fairness, whereas we focus on fairness among groups of clients where the goal is to create equitable learning scenarios among different groups. A recent study [26] also focuses on ensuring fair performance for both groups of clients and individual clients. However, prior knowledge of the client groups is required, while our approach automatically clusters the clients based on the activation vectors. Another concurrent work [27] orthogonal to ours uses the bias’s gradients in the neural network’s last layer to construct clusters and then perform heterogeneity-aware client sampling. This work poses an immense computational overhead, requiring computing the similarity matrix amongst all clients to sample the participating clients. Other clustering-based works [28, 29] perform clustered federated learning to promote personalization, which is not the objective of our work.
@@ 4.2 Framework: Equitable-FL
We emphasize that while K-Means relies on specifying the number of clusters as a hyperparameter, it is not unique in this requirement. Many other clustering algorithms, e.g. hierarchical clustering, also depend on a thresholding mechanism to determine cluster formation, which similarly requires fine-tuning for optimal results. Thus, the need for parameter adjustment is a common aspect across clustering techniques.
@@ 4.3 Convergence Analysis
Suppose Assumptions 1, 2, 3, and 4 hold true for Equitable-FL (refer Algorithm 1). In Equitable-FL set $\eta=\frac{1}{4E\sqrt{3LK}}$. Define a distribution $\mathbb{P}$ for $k\in\{0,\ldots,K-1\}$ such that $\mathbb{P}(k)=\frac{(1+\zeta-1)^{(K-1-k)}}{\sum_{k=0}^{K-1}(1+\zeta)^{k}}$ where $\zeta:=\eta^{2}E^{2}\left(9\eta L^{2}E+4\eta\mu E\right.\\ \left.+6L\left(1+\frac{4\eta^{2}\mu^{2}E^{2}}{18}\right)\right)$. Sample $k^{*}$ from $\mathbb{P}$. Then for $\eta LE\leq\frac{1}{2}$, $\eta\mu E\leq\frac{1}{2}$, $\mu<1$, and $K\geq max\left(\frac{3L}{32},\frac{\mu^{2}}{12L},\frac{\mu^{2}}{108L^{3}}\right)$; we have:
$\mathbb{E}[\|\nabla f(\bm{w}_{k^{*}})\|^{2}]\leq\frac{16\sqrt{3L}f(\bm{w}_{0})}{(1-\mu)\sqrt{K}}+\frac{\sqrt{L}\sigma^{2}}{2\sqrt{3K}E(1-\mu)}+\frac{\sigma^{2}}{36E^{2}K(1-\mu)}+\frac{(2+E)L\sigma^{2}}{36K(1-\mu)}+\frac{\mu\sigma^{2}}{6ELK(1-\mu)}\\ +\frac{\sigma^{2}}{18CELK(1-\mu)}\left(L^{2}+\frac{\mu}{2}\right)\sum_{q=1}^{C}\frac{1}{|\gamma_{q}|}.$ (8)
and the expectation is with respect to the randomness in all stochastic gradients and the random selection of $k$ according to the distribution $\mathbb{P}$.
As stated above, the analysis is conducted without restrictive assumptions such as convexity and bounded client dissimilarity [5, 35]. Broadly speaking, we can observe that it consists of two terms in eq. 8. In particular, the first term captures the impact of initialization, and the second set of terms results from the noisy stochastic updates of the clients. So, by setting the $\eta$ as described in the theorem, we see that Equitable-FL enjoys a convergence of $\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)$ to reach an $\epsilon$-stationary solution for our setting.
@@ 6 Conclusion and Future Work
As previously mentioned, privacy is a cornerstone of Federated Learning (FL). However, activation vectors can lead to information leakage if the server is compromised. Despite this risk, privacy-preserving mechanisms exist to enable secure learning without compromising privacy [41, 42]. This paper aims to demonstrate the effectiveness of using activation vectors to cluster client groups and develop a fair weighting mechanism for these groups. We also propose that this clustering approach can be applied to personalized federated learning [43, 44]. This method effectively clusters client groups and ensures fair solutions for each group, maintaining fairness for all participants.
@@ Appendix A Theorems and Lemmas
We base our convergence analysis on a framework similar to that of [14]. However, our approach introduces distinct variations due to the addition of a proximal term in local updates and the equitable reweighting of these updates across clients, leading to differences in the analysis compared to [14].
@@
Now, note that the optimal step size will be $\eta=\mathcal{O}\left(\frac{1}{E\sqrt{LK}}\right)$. So, then let us pick $\eta=\frac{1}{4E\sqrt{3LK}}$. We need to have $\eta LE\leq\frac{1}{2}$ and $\eta\mu E\leq\frac{1}{2}$; which happens for $K\geq max\left(\frac{L}{12},\frac{\mu^{2}}{12L}\right)$. Furthermore, lets ensure that $\zeta K<\frac{1}{2}$; which happens for $K\geq max\left(\frac{3L}{32},\frac{\mu^{2}}{108L^{3}},\frac{\mu^{2}}{72L}\right)$. Therefore we should have $K\geq max\left(\frac{3L}{32},\frac{\mu^{2}}{12L},\frac{\mu^{2}}{108L^{3}}\right)$. So, plugging in $\eta=\frac{1}{4E\sqrt{3LK}}$ and $1-\zeta K\geq\frac{1}{2}$ in eq. 16; we get,
$\sum_{k=0}^{K-1}p_{k}\mathbb{E}[\|\nabla f(\bm{w}_{k})\|^{2}]\leq\frac{16\sqrt{3L}f(\bm{w}_{0})}{(1-\mu)\sqrt{K}}+\frac{\sqrt{L}\sigma^{2}}{2\sqrt{3K}E(1-\mu)}\\ +\frac{\sigma^{2}}{36E^{2}K(1-\mu)}+\frac{(2+E)L\sigma^{2}}{36K(1-\mu)}+\frac{\mu\sigma^{2}}{6ELK(1-\mu)}\\ +\frac{\sigma^{2}}{18CELK(1-\mu)}\left(L^{2}+\frac{\mu}{2}\right)\sum_{q=1}^{C}\frac{1}{|\gamma_{q}|}.$ (17)
This finishes the proof. ∎